OpenStudy (anonymous):

Find the area of a circle circumscribed about an equilateral triangle whose side is 18 inches long.

OpenStudy (anonymous):



OpenStudy (anonymous):

you'll note that it splits the angles in two. We can solve for the radius using trig. We have a 30-120-30 triangle: sin(30)=r/18 18sin(30)=r=9 We can now solve for the circle are: Area=πr2=π(9)2=81π=

OpenStudy (anonymous):

That last part should say = 254.47

OpenStudy (cwrw238):

you can find the area of each segment using A = 0.5 r^2 A - area of triangle where A = pi/3 radians area of triangle = 0.5*18^2 sin pi/3

OpenStudy (anonymous):

thx guys appreciated it :)

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):